# RECAST: Reparameterized, Compact weight Adaptation for Sequential Tasks

### v1

@@ Abstract
Incremental learning aims to adapt to new sets of categories over time with minimal computational overhead. Prior work often addresses this task by training efficient task-specific adapters that modify frozen layer weights or features to capture relevant information without affecting predictions on any previously learned categories. While these adapters are generally more efficient than finetuning the entire network, they still can require tens or hundreds of thousands of task-specific trainable parameters even for relatively small networks, making it challenging to operate on resource-constrained environments with high communication costs like edge devices or mobile phones. Thus, we propose Reparameterized, Compact weight Adaptation for Sequential Tasks (RECAST), a novel method that dramatically reduces the number of task-specific trainable parameters to fewer than 50 – several orders of magnitude less than competing methods like LoRA. RECAST accomplishes this efficiency by learning to decompose layer weights into a soft parameter-sharing framework consisting of a set of shared weight templates and very few module-specific scaling factor coefficients. This soft parameter-sharing framework allows for effective task-wise reparameterization by tuning only these coefficients while keeping the templates frozen. A key innovation of RECAST is the novel weight reconstruction pipeline called Neural Mimicry, which eliminates the need for training in our framework from scratch. Extensive experiments across six diverse datasets demonstrate RECAST outperforms the state-of-the-art by up to $3\%$ across various scales, architectures, and parameter spaces. Moreover, we show that RECAST’s architecture-agnostic nature allows for seamless integration with existing methods, further boosting performance.
@@ 1 Introduction
Incremental Learning (IL) aims to learn from a continuous stream of data or tasks over time. Generally speaking, the goal of IL methods is to avoid catastrophic forgetting (McCloskey & Cohen, 1989; French, 1999) while minimizing computational overhead. As summarized in Figure 1(a), prior work in IL can be separated into three themes. First, Rehearsal methods that learn what samples to retain to use on subsequent iterations to ensure their task-specific knowledge is retained (Caccia et al., 2020; Rebuffi et al., 2017; Chaudhry et al., 2019; 2021; Shin et al., 2017), but they can require large amounts of storage as the number of tasks grows. Second, Regularization methods that require no extra storage (Kirkpatrick et al., 2017; Kim et al., 2021; Aljundi et al., 2018; Nguyen et al., 2018; Zhang et al., 2020), but can result in network instability due, in part, to the plasticity-stability dilemma (Grossberg, 1982). Finally, third, Reconfiguration/Adaptation methods (Yoon et al., 2018; Rusu et al., 2022; Hung et al., 2019; Mallya & Lazebnik, 2018; Singh et al., 2020; Verma et al., 2021; Jin & Kim, 2022; Ge et al., 2023; Douillard et al., 2022; Wang et al., 2023) that learn task-specific parameters, but can result in large numbers of new parameters per-task that may not scale well. Even methods designed to be efficient, like those based on LoRA (Hu et al., 2021), still require tens or hundreds of thousands of task-specific trainable parameters for each task (e.g.(Wu et al., 2024; Wang et al., 2024; Zhang et al., 2024; Erkoç et al., 2023)) making them ill-suited for some applications in resource-constrained environments like mobile phones or edge devices.
@@
Figure 1: (a) Existing IL methods,i.e. Rehearsal, Regularization, Reconfiguration - exhibit various limitations in terms of model complexity, memory requirements, and training overheads. In comparison our proposed method, (b) RECAST can be used as a frozen backbone, allowing efficient reparameterization of any target module with a negligible number of parameter updates (order of $10^{-6}$) and accommodate any number of disjoint tasks.
@@
To address this, we propose Reparameterized, Compact weight Adaptation for Sequential Tasks (RECAST), which provides an extremely parameter-efficient approach for IL via reparameterization of a model’s pretrained weights. As illustrated in Figure 1(b), RECAST achieves its compact task-specific representation by finetuning only the coefficients for a set of shared weight templates, enabling us to reconfigure an entire network’s weights using fewer than 50 trainable parameters. This framework shares some principles with prior Template Mixing methods (e.g.(Plummer et al., 2022; Savarese & Maire, 2019)). However, these methods did not explore applications to incremental learning, or the reparameterization paradigms but rather evaluated their models on tasks like hierarchy discovery and parameter budget allocation. In addition, these methods required that a model is trained in the Template Mixing framework from scratch, making them unable to leverage large pre-trained models like DINO (Caron et al., 2021).
@@
To enable RECAST to use large pretrained models that would be prohibitively expensive to retrain in our desired framework, we propose Neural Mimicry, a weight reconstruction approach that learns a decomposition of a pretrained layer into our framework such that it emulates the original network. This enables us to leverage pretrained weights from any model with minimal additional processing (i.e., just a few minutes). Few works in the reparameterization theme of IL have explored reconstruction methods. Still, they focus primarily on generating weights for LoRA matrices typically through diffusion (Wu et al., 2024; Wang et al., 2024; Zhang et al., 2024; Erkoç et al., 2023) or used pretrained weights to initialize their LoRA components (Liu et al., 2024a; Si et al., 2024). Prior work learned additional predictor networks from scratch to generate weights for some subset of modules (Salimans & Kingma, 2016). In contrast to these methods, our work directly changes the backbone weights rather than learning an adapter like LoRA-based methods, while also being several orders of magnitude more parameter efficient.
@@
We evaluate RECAST’s effectiveness on diverse datasets using Task-incremental IL settings. We compare against $12$ baselines comprising state-of-the-art IL methods on both CNN and Transformer architectures, where RECAST reports up to a 3% gain over prior work. When combined with adapter methods with higher parameter budgets, RECAST still boosts them by up to 2%.
@@
  • RECAST’s fine-tuned resource-performance control suits efficient continual learning with bounded resource constraints. Its compact task-specific parameters enable easier distribution across large-scale systems and edge devices.
@@ 2 Reparameterized, Compact weight Adaptation for Sequential Tasks (RECAST)
Figure 2: RECAST decomposes pretrained weights into templates and coefficients. During transfer-learning the templates are kept frozen and only coefficients get tuned to generate module-specific task-wise reparameterized weights
@@
In Task Incremental Learning (TIL), a model sequentially learns D tasks $\{d_{1},d_{2},d_{3},\dots d_{D}\}$, where each task $d_{i}=(X_{i},Y_{i})$ introduces new categories, domains, or output spaces. The objective is to maximize classification performance across all tasks while minimizing resource usage, i.e., $\max\sum_{i=1}^{D}P(d_{i})\text{ s.t. }R\leq R_{max}$ where $P(d_{i})$ is the performance on task $d_{i}$, $R$ is the total resource usage, and $R_{max}$ is the resource constraint. This resource-awareness is crucial in practice, as it enables deployment on edge devices with limited memory, impacts cloud infrastructure costs and real-time processing, and ensures accessibility in resource-limited environments. Thus, TIL poses a complex optimization problem: how can we design a system that adapts to new tasks without significantly increasing its computational footprint or compromising performance on existing tasks?
We address this multifaceted challenge with our novel frameworkReparameterized, Compact weight Adaptation for Sequential Tasks (RECAST), that strategically balances model plasticity and stability while adhering to the resource constraints in TIL scenarios. As illustrated in figure 2 RECAST introduces a flexible way to adapt neural network layers through the dynamic combination of sets of shared weight blueprints (i.e., Template Banks) and module-wise calibration factors or Coefficients. Our work builds upon and extends several key areas in neural network adaptation. In (Section 2.1) we present our Weight Decomposition scheme. It has been previously demonstrated that such modular networks are particularly suited for task adaptation (Mendez & Eaton, 2021; Ostapenko et al., 2021). Older works like, Weight Normalization (Salimans & Kingma, 2016) decouples the length of target weight vectors from their direction to reparameterize the network weights. Decomposition has also been used to make DNNs more interpretable Zhou et al. (2018) and to discover hierarchical structures Patil et al. (2023). Works by  Savarese & Maire (2019) and  Plummer et al. (2022) demonstrate the effectiveness of weight sharing through decomposition, similar to our approach. However, these methods either require training models from scratch in their decomposed form or learning additional task-driven priors to accommodate new tasks Veniat et al. (2021). RECAST introduces Neural Mimicry to enable the direct reconstruction of pretrained weights into reusable templates and sparse coefficients (Section 2.2). Prior approaches to generating neural network weights have explored various strategies: HyperNetwork (Ha et al., 2017) uses smaller networks to predict weights of larger networks. However, while we use static templates to reconstruct weights, HyperNetworks learns to predict weights from input vectors and require end-to-end training. NeRN (Ashkenazi et al., 2023) builds upon this idea of subnetworks by using positional embeddings of weight coordinates, but like HyperNetworks, doesn’t leverage pretrained knowledge. A recent work, DoRA (Liu et al., 2024b) shares some similarity with our approach by decomposing pretrained weights into magnitude vectors and LoRA matrices. However, in their case, the weight is utilized only during initialization and requires significant parameter tuning. Another contemporary work, Wang et al. (2024) attempts novel weight generation through autoencoders and latent diffusion. Again, their approach requires complex training of generative models. In contrast, RECAST offers a direct and efficient approach to simple decomposition (Algorithm  1) and novel weight generation, formally described in Algorithm 2. Our simple implementation allows the pipeline to finish executing in $<5$ minutes in our experiments, making the added overhead negligible (See Appendix Table 5). These two techniques also combine to establish RECAST as a parameter-efficient model adaptation scheme like LoRA (Chen et al., 2022; Liang & Li, 2024; Zhu et al., 2024; Nikdan et al., 2024) which is now popularly used to adapt to new tasks through low-rank updates. In fact, in Section 3, we empirically show, how RECAST can enhance the variants of LoRA providing a substantial boost in performance across a range of parameters.
@@ 2.1 Weight Decomposition
7:$M_{l}$: Number of modules in layer $l$
@@
14:              $g=\lceil l/(L/G)\rceil$triangleright Map layer to group
15:              $C_{l,m}={C_{l,m,i}^{k}\in\mathbb{R}|k=1,\ldots,K;i=1,\ldots,n}$triangleright Module coefficients
@@
RECAST maintains overall architecture (ResNet, Vision Transformer etc.) while replacing specific modules with components that can dynamically generate weights. A target module has access to a bank of templates which is shared with some other layers according to a specific grouping scheme (See Section 3.2). Each module has its own set of coefficients which are used to generate the weights. The process can be mathematically formalized as follows: Let $L$ be the total number of layers in the network organized into $G$ groups, where each group shares a template bank. For each group $g\in{1,\ldots,G}$, we define a template bank $\tau_{g}={T_{g,1},T_{g,2},\ldots,T_{g,n}}$, where $n$ is the number of templates per bank. Each template $T_{g,i}\in\mathbb{R}^{d_{out}\times d_{in}\times\ldots}$ has the same shape as the layer being parameterized. For each target module $m\in{1,\ldots,M_{l}}$ in layer $l\in{1,\ldots,L}$, where $M_{l}$ is the number of target modules in layer $l$, we generate $K$ sets of coefficient vectors. The $k$-th set of coefficients for module $m$ in layer $l$ is defined as $C_{l,m}^{k}={C_{l,m,1}^{k},C_{l,m,2}^{k},\ldots,C_{l,m,n}^{k}}$, where $C_{l,m,i}^{k}\in\mathbb{R}$ is the learned coefficient for template $T_{g,i}$ in group $g=\lceil l/(L/G)\rceil$. The weight matrix for module $m$ in layer $l$ is computed via:
@@
Here, $\cdot$ represents element-wise multiplication. The final weight matrix for module $m$ in layer $l$ is the average of these $K$ parameter matrices.
@@
2.2 Neural Mimicry
We introduce ”Neural Mimicry” as a foundational technique for efficiently emulating complex pretrained neural networks. This end-to-end pipeline enables the recreation of a target model within our framework, bypassing resource-intensive pretraining. The input to this pipeline is a pretrained target model $M$, whose module weights we aim to emulate. The output is a set of template banks and coefficients that, when combined, approximate the target module weights of $M$. The core of Neural Mimicry lies in its iterative refinement process, which progressively shapes the custom model $M^{*}$ to closely mirror the behavior of the target model $M$. The pretrained weights of M are obtained from official timm repository (See Appendix A.3). To generate weights for each module in each layer, we employ Equations 1 and 2 as described in Section 2.1, and then aim to reduce the difference between the original pretrained weight ($W_{l,m}$) and RECAST-generated weight ($W_{l,m}^{*}$). The main update step uses gradient descent (Algorithm 2 line 9) on the coefficients and templates ($\eta$ is the learning rate): $C_{l,m,i}^{k}\leftarrow C_{l,m,i}^{k}-\eta\frac{\partial\mathcal{L}}{\partial C_{l,m,i}^{k}};T_{g(l),i}\leftarrow T_{g(l),i}-\eta\frac{\partial\mathcal{L}}{\partial T_{g(l),i}}$. The process can be formulated as the following optimization problem:
where $\epsilon=(\epsilon_{1},\ldots,\epsilon_{K})$ is a set of K-noise vectors sampled from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$, where $\epsilon_{i}\in\mathbb{R}^{n\times 1\times\dots}$. The noise vectors are added with the coefficients only during the reconstruction training process. Here, $L$ denotes total layers, $M_{l}$ target modules in layer $l$, $n$ templates per bank, and $\mathcal{L}$ the weight discrepancy loss. The choice of $\mathcal{L}$ is an important design decision(see Section 3.2). We found Smooth L1 loss most effective, balancing L1 and L2 properties. It offers smooth gradients for small differences and stability for larger errors, crucial for diverse weights across layers and stable fine-tuning performance (Figures 6(a)3). This outperforms L2/MSE (sensitive to outliers) and KL divergence (computationally heavy) (Hinton et al., 2015; Ashkenazi et al., 2023). After reconstruction, we calculate the layerwise feature similarity between $M$ and $M^{*}$ through Cosine Similarity metric to evaluate the quality(See Table 6).

### v2

@@ Abstract
Incremental learning aims to adapt to new sets of categories over time with minimal computational overhead. Prior work often trains efficient task-specific adapters that modify frozen layer weights or features to capture relevant information without affecting predictions on any previously learned categories. While these adapters are generally more efficient than finetuning the entire network, they still can require tens or hundreds of thousands of task-specific trainable parameters, even for relatively small networks. This is can be problematic for resource-constrained environments with high communication costs, such as edge devices or mobile phones. Thus, we propose Reparameterized, Compact weight Adaptation for Sequential Tasks (RECAST), a novel method that dramatically reduces the number of task-specific trainable parameters to fewer than 50 – several orders of magnitude less than competing methods like LoRA. RECAST accomplishes this efficiency by learning to decompose layer weights into a soft parameter-sharing framework consisting of a set of shared weight templates and very few module-specific scaling factor coefficients. This soft parameter-sharing framework allows for effective task-wise reparameterization by tuning only these coefficients while keeping the templates frozen. A key innovation of RECAST is the novel weight reconstruction pipeline called Neural Mimicry, which eliminates the need for training in our framework from scratch. Extensive experiments across six diverse datasets demonstrate RECAST outperforms the state-of-the-art by up to $\sim 1.5\%$ and improves baselines $>3\%$ across various scales, architectures, and parameter spaces. Moreover, we show that RECAST’s architecture-agnostic nature allows for seamless integration with existing methods, further boosting performance [^1] Code: Repository.
@@ 1 Introduction
Incremental Learning (IL) aims to learn from a continuous stream of data or tasks over time. Generally speaking, the goal of IL methods is to avoid catastrophic forgetting (McCloskey & Cohen, 1989; French, 1999) while minimizing computational overhead. As summarized in Figure 1(a), prior work in IL can be separated into three themes. First, Rehearsal methods that learn what samples to retain to use on subsequent iterations to ensure their task-specific knowledge is retained (Caccia et al., 2020; Rebuffi et al., 2017; Chaudhry et al., 2019; 2021; Shin et al., 2017), but they can require large amounts of storage as the number of tasks grows. Second, Regularization methods that require no extra storage (Kirkpatrick et al., 2017; Kim et al., 2021; Aljundi et al., 2018; Nguyen et al., 2018; Zhang et al., 2020), but can result in network instability due, in part, to the plasticity-stability dilemma (Grossberg, 1982). Finally, third, Reconfiguration/Adaptation methods (Yoon et al., 2018; Rusu et al., 2006; Hung et al., 2019; Mallya et al., 2018; Singh et al., 2020; Verma et al., 2021; Jin & Kim, 2022; Ge et al., 2023; Douillard et al., 2022; Wang et al., 2022a) that learn task-specific parameters, but can result in large numbers of new parameters per-task that may not scale well. Even methods designed to be efficient, like those based on LoRA (Hu et al., 2022), still require tens or hundreds of thousands of task-specific trainable parameters for each task (e.g.(Wang et al., 2024; Zhang et al., 2024; Erkoç et al., 2023)) making them ill-suited for some applications in resource-constrained environments like mobile phones or edge devices.
@@
Figure 1: (a) Existing IL methods, i.e. Rehearsal, Regularization, Reconfiguration - exhibit various limitations in terms of model complexity, memory requirements, and training overheads. In comparison our proposed method, (b) RECAST can be uses as a frozen backbone, allowing efficient reparameterization of any target module with a negligible number of parameter updates (order of $10^{-6}$) and can accommodate any number of disjoint tasks.
@@
To address this, we propose Reparameterized, Compact weight Adaptation for Sequential Tasks (RECAST), which provides an extremely parameter-efficient approach for IL via reparameterization of a model’s pretrained weights. As illustrated in Figure 1(b), RECAST achieves its compact task-specific representation by finetuning only the coefficients for a set of shared weight templates, enabling us to reconfigure an entire network’s weights using fewer than 50 trainable parameters. RECAST framework shares some principles with prior Template Mixing methods (e.g.(Plummer et al., 2022; Savarese & Maire, 2019)). However, these methods did not explore incremental learning or reparameterization, focusing instead on hierarchy discovery and parameter allocation. They also required training from scratch, preventing the use of existing pre-trained models.
@@
In contrast, our proposed Neural Mimicry approach allows RECAST to utilize existing large pretrained weights without costly retraining. This weight reconstruction approach learns to decompose pretrained layers into our framework, effectively emulating the original network’s weights. This enables us to leverage any pretrained model with minimal processing (often just a few minutes). Thus, REACAST addresses a gap in reparameterization schemes and IL research where reconstruction methods remain underexplored where recent work has focused on generating weights for LoRA, typically through diffusion (Wang et al., 2024; Zhang et al., 2024; Erkoç et al., 2023) or used pretrained weights to initialize their LoRA components (Liu et al., 2024a; Si et al., 2024). Some prior work also learned additional predictor networks from scratch to generate weights for a subset of modules (Salimans & Kingma, 2016). In contrast to these methods, our work directly changes the backbone weights rather than learning an adapter like LoRA-based methods, while also being several orders of magnitude more parameter efficient. Figure 2 provides an overview of our approach.
@@
We evaluate RECAST’s effectiveness on diverse datasets in a Task-incremental IL setting. We compare against $16$ baselines comprising state-of-the-art IL methods on both CNN and Transformer architectures, where RECAST reports $>3\%$ gain over prior work. When combined with adapter methods with higher parameter budgets, RECAST further increases performance by $\sim{1.5\%}$.
@@
  • RECAST’s fine-tuned resource-performance control, suits efficient continual learning with bounded resource constraints. Its compact task-specific parameters enable easier distribution across large-scale systems and edge devices.
2 Related Work
Our work builds upon and extends several key areas in neural network adaptation, in particular the area of weight decomposition. These methods have previously shown that such modular networks are well suited for task adaptation (Mendez & Eaton, 2021; Ostapenko et al., 2021). Older works like Weight Normalization (Salimans & Kingma, 2016) decouples the length of target weight vectors from their direction to reparameterize the network weights. Decomposition has also been used to make DNNs more interpretable (Zhou et al., 2018) and to discover hierarchical structures (Patil et al., 2023). Savarese & Maire (2019) and Plummer et al. (2022) demonstrate the effectiveness of weight sharing through decomposition, similar to our approach. However, these methods either require training models from scratch in their decomposed form or learning additional task-driven priors to accommodate new tasks (Veniat et al., 2020). In terms of parameter-efficient tuning, LoRA is now popularly used to adapt to new tasks through low-rank updates (Chen et al., 2022; Liang & Li, 2024; Zhu et al., 2024; Nikdan et al., 2024), but adds thousands of tunable parameters. In contrast, our approach learns to decompose the weights of a pretrained network to enable quick adaptation to new tasks with $<0.0002\%$ task parameters.
Our work is also similar to works that generate neural network weights. Earlier works like, HyperNetworks (Ha et al., 2017) uses a small subnetwork to predict the weights of a larger network. Unlike our reconstruction scheme, this requires end-to-end training on input vectors. NeRN (Ashkenazi et al., 2023) builds upon this idea of subnetworks by using positional embeddings of weight coordinates; however, it does not leverage pre-trained knowledge. A recent work, DoRA (Liu et al., 2024b) shares some similarities with our approach by decomposing pretrained weights into magnitude vectors and LoRA matrices. However, the weight is utilized only during initialization and requires significant parameter tuning. Another contemporary work, Wang et al. (2024) attempts novel weight generation through autoencoders and latent diffusion, which can be resource-heavy to train. In comparison, RECAST offers a direct, efficient approach to weight generation and reparameterization. We also show that RECAST can be combined with existing methods to boost performance.
@@ 3 Reparameterized, Compact weight Adaptation for Sequential Tasks (RECAST)
Figure 2: RECAST decomposes module weights into templates and coefficients, which are learned through reconstruction from pretrained weights (Section 3.2). These components are linearly combined to dynamically generate weights for a target layer (Section 3.1). During TIL training, the templates are kept frozen, and only task-specific coefficients learned to create new layer weights.
@@
In Task Incremental Learning (TIL) a model sequentially learns D tasks $\{d_{1},d_{2},d_{3},\dots d_{D}\}$, where each task $d_{i}=(X_{i},Y_{i})$ introduces new categories, domains, or output spaces. The objective is to maximize classification performance across all tasks while minimizing resource usage, i.e., $\max\sum_{i=1}^{D}P(d_{i})\text{ s.t. }R\leq R_{max}$, where $P(d_{i})$ is the performance on task $d_{i}$, $R$ is the total resource usage, and $R_{max}$ is the resource constraint. This resource-awareness is crucial in practice, as it enables deployment on edge devices with limited memory, impacts cloud infrastructure costs and real-time processing, and ensures accessibility in resource-limited environments. Thus, TIL poses a complex optimization problem: how can we design a system that adapts to new tasks without significantly increasing its computational footprint or compromising performance on existing tasks?
We address this multifaceted challenge with our novel framework Reparameterized, Compact weight Adaptation for Sequential Tasks (RECAST), that strategically balances model plasticity and stability while adhering to resource constraints. As illustrated in Figure 2, RECAST introduces a flexible way to adapt the layers of neural networks through the dynamic combination of sets of shared weight blueprints (i.e., template banks) and module-wise calibration factors or coefficients. In Section 3.1 we present our Weight Decomposition scheme. Section 3.2 introduces Neural Mimicry for direct reconstruction of pretrained weights into reusable templates and coefficients.
@@ 3.1 Weight Decomposition
7:$M_{l}$: Number of target modules in layer $l$
@@
14:              $C_{l,m}={C_{l,m,i}^{k}\in\mathbb{R}|k=1,\ldots,K;i=1,\ldots,n}$triangleright Module coefficients
[…]
@@
At its core RECAST replaces pretrained layer weights with modules that can dynamically generate weights. A target module has access to a bank of templates which is shared with some other layers in the network (discussed in Section 4.2). Each module has its own set of coefficients which are used to generate the weights. Formally, let $L$ be the total number of layers in the network organized into $G$ groups, where each group shares a template bank. For each group $g\in{1,\ldots,G}$, we define a template bank $\tau_{g}={T_{g,1},T_{g,2},\ldots,T_{g,n}}$, where $n$ is the number of templates per bank. Each template $T_{g,i}\in\mathbb{R}^{d_{out}\times d_{in}\times\ldots}$ has the same dimensions as the layer being parameterized. For each target module $m\in{1,\ldots,M_{l}}$ in layer $l\in{1,\ldots,L}$, where $M_{l}$ is the number of target modules in layer $l$, we generate $K$ sets of coefficient vectors. The $k$-th set of coefficients for module $m$ in layer $l$ is defined as $C_{l,m}^{k}={C_{l,m,1}^{k},C_{l,m,2}^{k},\ldots,C_{l,m,n}^{k}}$, where $C_{l,m,i}^{k}\in\mathbb{R}$ is the learned coefficient for template $T_{g,i}$ in group $g=\lceil l/(L/G)\rceil$. The weight matrix for module $m$ in layer $l$ is computed via:
@@
where $\cdot$ represents element-wise multiplication. The final weight matrix for module $m$ in layer $l$ is generated by averaging these $K$ parameter matrices.
@@
3.2 Neural Mimicry
We introduce “Neural Mimicry” as a foundational technique for efficiently emulating complex pretrained neural networks. This end-to-end pipeline enables the recreation of a target model within our framework, bypassing resource-intensive pretraining. The input to this pipeline is a pretrained target model $M$, whose module weights we aim to emulate. The output is a set of template banks and coefficients that, when combined, approximate the target module weights of $M$. The core of Neural Mimicry lies in its iterative refinement process, which progressively shapes the custom model $M^{*}$ to closely mirror the behavior of the target model $M$. In our experiments the pretrained weights of M are obtained from official releases (see Appendix A.3). To generate weights for each module in each layer, we employ Equations 1 and 2 as described in Section 3.1, and then aim to reduce the difference between the original pretrained weight ($W_{l,m}$) and RECAST-generated weight ($W_{l,m}^{*}$). The main update step uses gradient descent (Algorithm 2 line 9) on the coefficients and templates ($\eta$ is the learning rate): $C_{l,m,i}^{k}\leftarrow C_{l,m,i}^{k}-\eta\frac{\partial\mathcal{L}}{\partial C_{l,m,i}^{k}};T_{g(l),i}\leftarrow T_{g(l),i}-\eta\frac{\partial\mathcal{L}}{\partial T_{g(l),i}}$. The process can be formulated as the following optimization problem:
where $\epsilon=(\epsilon_{1},\ldots,\epsilon_{K})$ is a set of K-noise vectors sampled from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$, where $\epsilon_{i}\in\mathbb{R}^{n\times 1\times\dots}$. The noise vectors are added with the coefficients only during the reconstruction training process. Here, $L$ denotes total layers, $M_{l}$ target modules in layer $l$, $n$ templates per bank, and $\mathcal{L}$ the weight discrepancy loss. As discussed in Section 4.2, $\mathcal{L}$ should be chosen carefully, and we found Smooth L1 loss to be most effective for generating diversified template. It offers smooth gradients for small differences and stability for larger errors, crucial for stable reconstruction (Figures 5 & 4). This outperforms L2/MSE (sensitive to outliers) and KL divergence (computationally expensive) (See Figures 5 &  6). After reconstruction, we calculate the layerwise feature similarity between $M$ and $M^{*}$ using the cosine similarity metric to evaluate the quality (see Table 6). Overall, this process requires negligible overhead ($<5$ minutes, see Table 5).